OpenStudy (anonymous):

Divide and express the result in standard form: 8i/(4-3i)

OpenStudy (lgbasallote):

\[\frac{8i}{4-3i} \times \frac{4+3i}{4+3i}\] \[\frac{8i(4+3i)}{16 - 9i^2}\] \[\frac{8i(4+3i)}{16 - 9(-1)}\] \[\frac{8i(4+3i)}{16 + 9}\] \[\frac{8i(4+3i)}{25}\] simplify by distribution